Circular No. 6495

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 Q1 (TABUR) K. Dennerl, J. Englhauser, and J. Trümper, Max-Planck-Institut für

Extraterrestrische Physik, Garching; and C. Lisse, Goddard Space Flight Center, NASA, communicate: “X-ray emission from comet C/1996 Q1 was detected with the ROSAT High Resolution Imager. The observations were performed on several occasions between Sept. 26 and Oct. 6, when Δ = 0.51–0.42 AU and r = 1.09–0.99 AU. The x-ray emission was highly variable on time scales of hours (or less), both in intensity (at least by a factor of five) and morphology, with episodes of considerable deviation from the crescent shape of C/1996 B2 (Hyakutake) observed under similar phase angles (IAUC 6373). The brightest parts of the emission were always concentrated in an area with a typical extent of 5′ (90 000 km) and displaced sunward from the nucleus by ∼ 100′′ (30 000 km). They reached a peak surface brightness of 0.01 count s^{−1} arcmin^{−2}, similar to that of comet C/1996 B2 during its close approach to the earth in Mar. The rapid morphological changes seem to rule out any scattering of solar x-rays by material in the comet’s coma as the dominant mechanism for producing the bright parts of the x-ray emission.”

COMET P/1996 N2 (ELST-PIZARRO) H. Böhnhardt, University of Munich, reports on behalf of the European

Southern Observatory (ESO) Rosetta/ISO Target Comets Imaging Team: “Images of the comet were taken by R. Schulz (Max-Planck-Institut für Aeronomie, Katlenburg-Lindau), G. P. Tozzi (Arcetri Observatory), and H. Rauer (Meudon Observatory) on Sept. 16 and Oct. 4 with, respectively, the 2.2-m MPG/ESO telescope (+ EFOSC2) and the Danish 1.54-m telescope (+ DFOSC) at La Silla. These images show a long antitail in p.a. 254^{o}– 255^{o} and a previously undetected antisolar tail in p.a. 69^{o} (as predicted on IAUC 6473). Preliminary analysis by Z. Sekanina and myself indicates that dust was being released from the object over a period of many weeks or months until approximately mid-July 1996, when the comet was discovered (IAUC 6456). This evidence eliminates the possibility that the features are products of an instant event (e.g., a collision). Dust particles subjected to solar-radiation-pressure accelerations of up to 0.05 the solar attraction are clearly apparent in the antisolar tail.”

1996 October 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green